An advertising agency for a pineapple drink claims that the drink contains 15% pineapple juice. How much pure pineapple juice would have to be added to 8 quarts of the drink to obtain a mixture containing 50% pineapple juice?
Answer by checkley77(12569)

.15*8+X=.5(8+X)
1.2+X=4+.5X
X-.5X=4-1.2
.5X=2.8
X=2.8/.5
X=5.6 QUARTS OF PURE PINEAPPLE IS NEEDED.
PROOF:
.15*8+5.6=.5(8+5.6)
1.2+5.6=.5*13.6
6.8=6.8